# Improved Bag-of-Words Image Retrieval with Geometric Constraints for Ground Texture Localization

### v1

@@ I Introduction
Recognizing the importance of BoW to high-accuracy localization and its already prevalent use for ground texture localization, here we present several techniques to substantially improve BoW for this application. We leverage extensive previous work on BoW regarding approximate $k$-means (AKM) vocabularies and soft assignment, where descriptors can be assigned to several visual words, to greatly improve the effectiveness of the visual vocabulary. Furthermore, we exploit the consistent orientation and constant scale constraints inherent to ground texture localization to improve both precision and recall. Acknowledging the differing requirements of global localization and loop closure in SLAM, we propose two tailored versions of our algorithm: a high-accuracy version (localization frequency $>5~{}\text{Hz}$) for global localization and a higher-speed version ($>13~{}\text{Hz}$) optimized for loop closure detection for SLAM systems. Notably, this work can be immediately applied to previous ground texture localization works that use BoW to directly improve their accuracy and overall results. To facilitate this, we have released our code on our lab website.
@@ III-A2 General Timing
[…]
@@ III-C Loop Closure Detection
Finally, we evaluate the loop closure detection performance of our algorithms. As a BoW query is often the first step in a loop closure pipeline before additional filtering, a useful evaluation metric is Recall@$N$. Recall@$N$ measures the percentage of queries that have at least one correct result when the top $N$ images are returned [32]. A correct match is defined as a returned image that overlaps the query image by at least 25%. The Recall@$N$ provides insight into the algorithm’s ability to retrieve any true positives before subsequent filtering steps.
Fig. 5 (Top Left) plots the Recall@$N$ for DBoW and our algorithms across all asphalt test set images. Both of our algorithms significantly outperform DBoW in Recall@$N$, with our high-accuracy algorithm achieving near-perfect Recall@$N$ and demonstrating exceptional loop closure detection capability.
Figure 5: Top Left: The Recall@$N$ for DBoW and our algorithms on the asphalt dataset. Other: BoW similarity scores for potential loop closures for DBoW and our algorithms.
While Recall@$N$ indicates an algorithm’s ability to retrieve potential loop closures, it does not reflect the distinguishability of true positives from false positives. Typically, SLAM systems using BoW employ a filtering step, rejecting images with BoW similarity scores below a defined threshold since they are unlikely to be true loop closures. Ideally, a large difference exists between the scores of true and false closures to improve loop closure recall and consequently localization accuracy.
@@ IV Conclusion
[…]

### v2

@@ I Introduction
Recognizing the importance of BoW to high-accuracy localization and its already prevalent use for ground texture localization, here we present several techniques to substantially improve BoW for this application. We leverage extensive previous work on BoW regarding approximate $k$-means (AKM) vocabularies and soft assignment, where descriptors can be assigned to several visual words, to greatly improve the effectiveness of the visual vocabulary. Furthermore, we exploit the consistent orientation and constant scale constraints inherent to ground texture localization to improve both precision and recall. Acknowledging the differing requirements of global localization and loop closure in SLAM, we propose two tailored versions of our algorithm: a high-accuracy version (localization frequency $5~\text{Hz}$) for global localization and a higher-speed version ($13~\text{Hz}$) optimized for loop closure detection for SLAM systems. Notably, this work can be immediately applied to previous ground texture localization works that use BoW to directly improve their accuracy and overall results. To facilitate this, we have released our code on our lab website.
@@ III-A2 General Timing
DBoW requires 6.7 ms to transform ORB features into BoW vectors, which is faster than the 24 ms needed by our algorithms. This difference is attributed to DBoW’s use of an HKM vocabulary, in contrast to our use of an AKM vocabulary. DBoW is also the fastest for database insertion, but given that all times are below 2 ms, this difference is negligible for overall runtime. Our high-speed algorithm version without soft assignment achieves the fastest query time since the size binning makes the BoW vectors larger and consequently results in a sparser inverse index. Meanwhile, our high-accuracy algorithm with soft assignment is the slowest since the additional word assignments make the inverse index more dense. However, with a query frequency exceeding $5~\text{Hz}$, we feel that this tradeoff for the significant accuracy improvement shown in the previous section is justifiable.
@@ III-C Loop Closure Detection
Finally, we evaluate the loop closure detection performance of our algorithms. As a BoW query is often the first step in a loop closure pipeline before additional filtering, a useful evaluation metric is Recall@NN. Recall@NN measures the percentage of queries that have at least one correct result when the top NN images are returned [32]. A correct match is defined as a returned image that overlaps the query image by at least 25%. The Recall@NN provides insight into the algorithm’s ability to retrieve any true positives before subsequent filtering steps.
Fig. 5 (Top Left) plots the Recall@NN for DBoW and our algorithms across all asphalt test set images. Both of our algorithms significantly outperform DBoW in Recall@NN, with our high-accuracy algorithm achieving near-perfect Recall@NN and demonstrating exceptional loop closure detection capability.
Figure 5: Top Left: The Recall@NN for DBoW and our algorithms on the asphalt dataset. Other: BoW similarity scores for potential loop closures for DBoW and our algorithms.
While Recall@NN indicates an algorithm’s ability to retrieve potential loop closures, it does not reflect the distinguishability of true positives from false positives. Typically, SLAM systems using BoW employ a filtering step, rejecting images with BoW similarity scores below a defined threshold since they are unlikely to be true loop closures. Ideally, a large difference exists between the scores of true and false closures to improve loop closure recall and consequently localization accuracy.
@@ IV Conclusion
We have presented several modifications to the standard BoW algorithm that significantly enhance its performance for both global localization and loop closure detection in ground texture applications. These modifications include size binning, orientation verification, AKM vocabulary usage, and soft assignment. We introduce a high-accuracy version, leveraging soft assignment, that achieves state-of-the-art global localization at over $5~\text{Hz}$, and a high-speed version (without soft assignment) that provides robust localization at over $13~\text{Hz}$. Both versions are readily applicable to existing BoW-based localization and SLAM systems for ground textures, and their implementations are available on our lab website. Future work will focus on refining other aspects of the SLAM pipeline to further exploit the geometric constraints of ground texture localization, ultimately improving the overall SLAM capabilities of these systems.